OpenStudy (anonymous):

Help please..The table lists data regarding the average salaries of several professional athletes in the years 1991 and 2001. A)use the date points to find a linear function that fits the date. B) use the function to predict the average salary in 2005 and 2010. Year….Av erage Salary. 1991 $264,000 2001 $1,450,000 A linear function that fits the date is s(x)=___. (Let x=the number of years since 1990, and let s= the average salary years from 1990). The predicted average salary for 2005 is $___ The predicted average salary for 2010 is $___ (round to the nearest whole number)

OpenStudy (amistre64):

M(y) is what I would have named it: Money (year)

OpenStudy (amistre64):

you are given 2 points; subtract one point from the other and stack M/y to determine a slope

OpenStudy (anonymous):

i got my first part, but stuck at how to get the 2005 prediction

OpenStudy (anonymous):

y=1,444,000. i think i got that right, just need to determine the answer for 2005 and2010

OpenStudy (amistre64):

2001 $1,450,000 -1991 - 264,000 ------------------- 10 $1,186,000 slope = M/y: 1,186,000/10 = 118,600 now we can use this slope to determine an equation that will give us predictions for a given year: M(y) = 1,186,000y - 2,361,062,000 this of course is given for the year format xxxx, as opposed to the "how far from" stuff

OpenStudy (amistre64):

M(2005) = whatever the calculate makes of that :)

OpenStudy (amistre64):

I get: $16,868,000 hopefully

OpenStudy (anonymous):

is that for 2005?

OpenStudy (amistre64):

thats what I get when I input "2005" into my equation; if you have already calibrated it for "years after" you simply have to make: x = 4 in your own equation and it should equal out

OpenStudy (amistre64):

err, 1991 to 2005 aint 4 years, had a synapse explode

OpenStudy (amistre64):

2,3,4,5,6,7,8,9,0,1,2,3,4,5 that would be 14 years after 1991

OpenStudy (amistre64):

try x=14 in your first equation and see what you get

OpenStudy (anonymous):

ok, hang on

OpenStudy (amistre64):

i see a typo in mine thats persisting; il correct it and see if we agree

OpenStudy (amistre64):

M(y) = 118,600y - 236,396,600 M(2005) = 118,600(2005) - 236,396,600 = 1,396,400 that should be better

OpenStudy (amistre64):

im gonna have to do this the long way .....

OpenStudy (anonymous):

how did you come up with the 236,396,600?

OpenStudy (amistre64):

I used the slope intercept form of the line equation and simplified

OpenStudy (amistre64):

or at least I tried to, but i think I got mixed up in it

OpenStudy (anonymous):

If you know how to get an equation of a line using two points, then do that - that way your independent variable is just the year instead of a number like 14

OpenStudy (anonymous):

\[\text{My}(y)=118600 y-2.35869\times 10^8 \]My(y) evaluated at 1991, 2001, 2005 and 2010 is:\[\{264000.,1.45\times 10^6,1.9244\times 10^6,2.5174\times 10^6\} \]

OpenStudy (amistre64):

(2001,1450000) we have one of the points: we have a slope: M(y) = my-mPy+PM M(y) = 118600y-118600Py+PM M(y) = 118600y-118600(2001)+(1450000) M(y) = 118600y-237318600+(1450000) M(y) = 118,600y-235,868,600 should work

OpenStudy (anonymous):

robtoebey - I'd pass that through N[#, 12] so it doesn't round

OpenStudy (amistre64):

that does the job :) just had to remember how to add and multiply stuff

OpenStudy (anonymous):

i'm confused...

OpenStudy (amistre64):

the big numbers are prolly whats confusing you

OpenStudy (anonymous):

yes.. just confusing how to figure for 2005 and 2010

OpenStudy (anonymous):

so i subtract 118,600y-235,868,600 to get my answer for 2005?

OpenStudy (amistre64):

given 2 points (x1,y1) and (x2,y2) we can find the slope,m, by subtracting the points and ratio then as y/x (x1,y1) - (x2,y2) ----------- x1-x2, y1-y2 m = \(\cfrac{y1-y2}{x1-x2}\) knowing m we can convert the slope formula into the point slope form \[\frac{y1-y2}{x1-x2}=m\] \[{y1-y2}=m({x1-x2})\] this in turn can be rearranged into slope intercept form: \[{y1-y2}=m({x1-x2})\] \[{y1}=m({x1-x2})+y2\] \[{y1}=mx1-mx2+y2\] \[{y1}=mx1+(-mx2+y2)\] \[{y1}=mx1+b\] make y1 and x1 generic \[{y}=mx+b\]

OpenStudy (amistre64):

you simpy use the formula you developed in part "a" to find the answer for part "b"

OpenStudy (amistre64):

since you say you did "a" already, what formula did you develop?

OpenStudy (anonymous):

OK, I'll save this formula and see if I can figure this out, I am just confused, first time taking algebra, doing online classes and now I have to figure how to do algebra 1, ugh..frustrating.

OpenStudy (amistre64):

once i determine the slope; i use any given point to finish up the equation for part "a" \[{y}=mx+(-mx2+y2)\] \[{y}=118,600x+(-118,600(1991)+(264,000))\] \[{y}=118,600x+(-235,868,600)\] \[{y}=118,600x-235,868,600\] when the year: x = 2005, we get \[{y}=118,600(2005)-235,868,600\] \[{y}=1,924,400\]

OpenStudy (anonymous):

part a answer. i came up with y=144,000 m=264,000-1,450,000/10 years=1,186,000 divided by 10=118,600 y=264,000+(10*118,600) y=144,000

OpenStudy (amistre64):

m is good

OpenStudy (amistre64):

the other parts are a bit odd to me

OpenStudy (anonymous):

i am not use to using the graphing tools on here, sorry

OpenStudy (anonymous):

oh ok..so i see what you are doing..

OpenStudy (amistre64):

good, cause I got no idea what you were doing with this: y=264,000+(10*118,600)

OpenStudy (anonymous):

haha..my apologies..

OpenStudy (anonymous):

i appreciate your help amistre, now i am going so see if i can figure for 2010 :) here's a medal for you help, thanks bunches..

OpenStudy (amistre64):

youre welcome, and good luck ;)

OpenStudy (anonymous):

is needed, thanks.

OpenStudy (anonymous):

ktklown "robtoebey - I'd pass that through N[#, 12] so it doesn't round" Refer to the Mathematica attachment. To effect the result you suggested requires the "Round" function in Mathematica.